40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 29 13:54:08 2008, rana, Update, SUS, ETMY and PRM have EQ related problems Untitled.png
    Reply  Tue Jul 29 14:38:02 2008, rob, Update, SUS, ETMY and PRM have EQ related problems 
       Reply  Tue Jul 29 19:53:19 2008, Koji, Update, SUS, PRM photos from the south window PRM_from_South_Window1.jpgPRM_from_South_Window2.jpg
Message ID: 755     Entry time: Tue Jul 29 13:54:08 2008     Reply to this: 756
Author: rana 
Type: Update 
Category: SUS 
Subject: ETMY and PRM have EQ related problems 
The attached trend shows that ETMY and PRM both had large steps in their sensors
around the time of the EQ and didn't return afterwards. The calibration of the
OSEM sensors is ~0.5 mm/V. The PRM sensors respond when we give it huge biases
but there is very little change in the ETMY. Almost certainly true that the
optics have shifted in their wire slings and that we will have to vent to
examine and repair at least ETMY.

Jenne is looking at the spectra of the other suspensions to see if there is
other more subtle issues.
Attachment 1: Untitled.png  19 kB  | Hide | Hide all
Untitled.png
ELOG V3.1.3-